Susie is 7 years old. She is 1 year older than one-eighth the age of her father. How old is Susie's father?
Answer by xoapplexx(2)

S:Susie's age
F:father's age
S=1+F/8
7=1+F/8
Subtract 1 from both sides.
6=F/8
times 8 on both sides.
48=F
F=48
The father's age is 48.